# MetaChest: Generalized few-shot learning of patologies from chest X-rays

### v1

@@ 6.2 Leveraging ImageNet
As shown in Table 4, BatchBased consistently outperforms ProtoNet-ML across all models and metrics. For instance, on ImageNet-1K, BatchBased surpasses ProtoNet-ML by 4.31 $HM$ points. When comparing the ImageNet-1K and ImageNet-21K models for BatchBased, the former achieves better results across all metrics. For example, ImageNet-1K yields an improvement of 0.78 $HM$ points compared to ImageNet-21K. Furthermore, BatchBased initialized with ImageNet-21K weights demonstrates a 4.47 $HM$ point gain over randomly initialized models. The literature on few-example regimes in inter-domain scenarios reports inconclusive findings regarding the benefits of using pre-trained models on ImageNet-1K Cherti & Jitsev [8]. However, our results indicate that using pre-trained models consistently improves performance on chest X-ray images.
@@ 6.3 Few-shot learning vs transfer learning
We observe that ProtoNet-ML achieves improved performance in only a limited subset of task configurations. Table 5 shows that these improvements occur primarily in the 1-unseen setting and tend to disappear as the number of shots ($k$-shot) or classes ($n$-way) increases. In the remaining task configurations, BatchBased outperforms ProtoNet-ML. Moreover, as shown in Figure 5, BatchBased demonstrates improved performance as the number of shots increases, whereas ProtoNet-ML’s performance remains nearly constant. These findings are consistent with those reported by Luo et al. [30], who compared the performance of conventional models with meta-learning methods such as ProtoNet on SFSC tasks across various natural image datasets. They found that conventional models tend to scale better than meta-learning approaches, particularly on fine-grained datasets. In medical applications, where datasets often include many classes and dozens of examples per class, our results suggest that BatchBased is the more effective approach for training pathology classifiers.
@@ 6.4 Pathology Classification Complexity
We observe that, as the number of classes in the episode increases while the number of unseen classes is held constant, performance improves consistently. For example, in Table 6, the 5-way, 3-unseen configuration with 15-shot outperforms the 4-way, 3-unseen configuration with the same number of shots by 0.99 $HM$ points, highlighting the performance gain from adding a single class. As shown in Figure 6, the 4-way configurations (solid orange line) consistently achieve higher $HM$ scores than the 3-way configurations (solid blue line), while the 5-way configurations (solid green line) outperform the 4-way configurations (solid orange line). This suggests that increasing the number of classes per episode, and consequently the number of examples, reduces task complexity and leads to improved performance. Similar results have been reported in the SFSC literature on natural image datasets, where a higher number of classes per episode consistently improves classification performance [30].
@@ 6.5 X-ray resolution
In this experiment, we trained models using three different architectures and systematically varied the X-ray resolution to study its impact. Images were resized using the Lanczos algorithm, a high-quality resampling method known for preserving edge sharpness and fine details [2]. The results are summarized in Table 7.
@@
These results are consistent with previous findings in the medical imaging literature under a complete data regime. In mammography, for example, lesions are detected more accurately in images with a resolution of $1700\times 2100$ pixels [42]. Similarly, in the case of chest X-rays, Rochmawanti & Utaminingrum [44] compared the performance of two models on the ChestX-ray14 dataset using resolutions of $64\times 64$ and $320\times 320$ pixels, finding better performance with higher resolution.
For MobileNetV3-Small-0.75, performance progressively improves as the resolution increases up to $768\times 768$, but begins to decline at higher resolutions, as shown in Figure 7. In contrast, both ConvNeXt-Tiny and DenseNet-121 exhibit consistent improvements with increasing resolution.
ConvNeXt-Tiny outperforms the other two architectures across all evaluated resolutions. The highest performance is achieved with this architecture at a resolution of $768\times 768$, although it surpasses the best result of MobileNetV3-Small-0.75 by only 1.8 $HM$ points. ConvNeXt-Tiny improves upon MobileNetV3-Small-0.75 at the default resolution used in this work ($384\times 384$) by just 2.26 $HM$ points. This finding is relevant because higher resolutions entail an exponential increase in memory requirements for storing and processing images. Due to resource limitations, experiments with ConvNeXt-Tiny at $1024\times 1024$ and DenseNet-121 at $768\times 768$ and $1024\times 1024$ could not be performed.
@@ 6.6 Architectures
We investigate how connectivity patterns and the number of parameters/operations influence pathology classification performance. This is particularly relevant because evidence from the complete data regime in language modeling [23, 4] and computer vision [26, 43] suggests that increasing network size and training data consistently reduces error. For meta-learning, Chen et al. [6] evaluated several convolutional architectures in few-shot multi-class classification on mini-ImageNet and a reduced version of the CUB dataset [56]. Their findings are inconclusive: while deeper architectures improved performance on CUB, gains on mini-ImageNet were observed only in certain cases.
In this experiment, we study the effect of convolutional and attention-based connectivity patterns using popular vision architectures. For both types, we examine efficient architectures with relatively few parameters and operations, as well as larger networks. For efficient convolutional architectures, we focus on ConvNeXt-Atto [58] and lightweight versions of MobileNet [16], while for Transformer-based models, we use MobileViTV2-1.0 [34]. For larger convolutional architectures, we experiment with DenseNet-121, DenseNet-161 [18], and ConvNeXt-Tiny, and for Transformer-based models, we consider MobileViTV2-2.0 [34]. Table 8 summarizes the results, comparing the performance of these architectures along with their number of parameters and operations.
@@
Among the larger architectures, ConvNeXt-Tiny achieves the highest performance, reaching 85.22 $HM$ points. Notably, it outperforms DenseNet-161 by 1.32 $HM$ points, an architecture previously shown to be effective for medical image analysis problems [44, 60, 31, 9].
@@
Interestingly, across both efficient and larger architectures, convolutional models outperform their Transformer-based counterparts. This trend is illustrated in Figure Figure 8, which depicts the relationship between model performance and computational efficiency for the evaluated architectures.
@@ 7 Conclusions and future work
The adoption of the generalized few-shot learning paradigm aligns more closely with the clinical presentation of pathologies in chest X-rays than the standard few-shot classification paradigm. This task formulation is particularly well-suited to specialized medical contexts and useful for a variety of scenarios within the healthcare domain. In addition, the proposed multi-label episode generation algorithm enables the creation of complex classification tasks, further broadening its applicability to real-world medical settings. Interestingly, our results show that BatchBased is an effective classification method in few-shot scenarios, despite being based on standard transfer learning and not specifically designed for few-shot learning. We also observed that increasing the number of classes per episode ($n$-way) and the number of training examples per class ($k_{trn}$) improves model performance by enhancing task robustness. With respect to image resolution, we found that using higher resolutions than those commonly applied in natural image tasks leads to better classification performance. This improvement is likely due to the fine-grained nature of pathology classification, where abnormal patterns are subtle and can be overlooked at lower resolutions. However, this performance gain comes at the cost of increased computational requirements and longer training times. Finally, we observed that employing efficient architectures can yield results comparable to those achieved by larger models, offering substantial advantages in terms of computational cost.
As future work, we envision three main research directions. First, leveraging Vision Foundation Models as a starting point for pathology classification. Pre-trained on large-scale datasets, these models could provide richer and more generalizable feature representations, thereby enhancing classification performance. Second, developing multimodal classification models that integrate complementary information from radiology reports, such as radiologist notes and clinical records. Incorporating this additional contextual information could enrich the diagnostic process and improve overall performance. Finally, conducting a comparative evaluation between model predictions and expert radiologist assessments. Such a study would enable clinical validation of the results and provide a more accurate and reliable measure of the model’s effectiveness in clinical settings.
@@ Acknowledgements
This work was supported by UNAM-PAPIIT grant IV100420. The first author was supported by the National Council of Science and Technology (CONACYT), Mexico, under scholarship number 710012.

### v2

@@ 6.2 Leveraging ImageNet
As shown in Table 4, BatchBased consistently outperforms ProtoNet-ML across all models and metrics. For instance, on ImageNet-1K, BatchBased surpasses ProtoNet-ML by 4.31 $HM$ points. When comparing the ImageNet-1K and ImageNet-21K models for BatchBased, the former achieves better results across all metrics. For example, ImageNet-1K yields an improvement of 0.78 $HM$ points compared to ImageNet-21K. Furthermore, BatchBased initialized with ImageNet-21K weights demonstrates a 4.47 $HM$ point gain over randomly initialized models. The literature on few-example regimes in inter-domain scenarios reports inconclusive findings regarding the benefits of using pre-trained models on ImageNet-1K cherti2022. However, our results indicate that using pre-trained models consistently improves performance on chest X-ray images.
@@ 6.3 Few-shot learning vs transfer learning
We observe that ProtoNet-ML achieves improved performance in only a limited subset of task configurations. Table 5 shows that these improvements occur primarily in the 1-unseen setting and tend to disappear as the number of shots ($k$-shot) or classes ($n$-way) increases. In the remaining task configurations, BatchBased outperforms ProtoNet-ML. Moreover, as shown in Figure 5, BatchBased demonstrates improved performance as the number of shots increases, whereas ProtoNet-ML’s performance remains nearly constant. These findings are consistent with those reported by luo2023closerlookagain, who compared the performance of conventional models with meta-learning methods such as ProtoNet on SFSC tasks across various natural image datasets. They found that conventional models tend to scale better than meta-learning approaches, particularly on fine-grained datasets. In medical applications, where datasets often include many classes and dozens of examples per class, our results suggest that BatchBased is the more effective approach for training pathology classifiers.
@@ 6.4 Pathology Classification Complexity
We observe that, as the number of classes in the episode increases while the number of unseen classes is held constant, performance improves consistently. For example, in Table 6, the 5-way, 3-unseen configuration with 15-shot outperforms the 4-way, 3-unseen configuration with the same number of shots by 0.99 $HM$ points, highlighting the performance gain from adding a single class. As shown in Figure 6, the 4-way configurations (solid orange line) consistently achieve higher $HM$ scores than the 3-way configurations (solid blue line), while the 5-way configurations (solid green line) outperform the 4-way configurations (solid orange line). This suggests that increasing the number of classes per episode, and consequently the number of examples, reduces task complexity and leads to improved performance. Similar results have been reported in the SFSC literature on natural image datasets, where a higher number of classes per episode consistently improves classification performance [luo2023closerlookagain].
@@ 6.5 X-ray resolution
In this experiment, we trained models using three different architectures and systematically varied the X-ray resolution to study its impact. Images were resized using the Lanczos algorithm, a high-quality resampling method known for preserving edge sharpness and fine details [interpolationalgorithms]. The results are summarized in Table 7.
@@
These results are consistent with previous findings in the medical imaging literature under a complete data regime. In mammography, for example, lesions are detected more accurately in images with a resolution of $1700\times 2100$ pixels [mammograms2018]. Similarly, in the case of chest X-rays, chestxrayresolutions compared the performance of two models on the ChestX-ray14 dataset using resolutions of $64\times 64$ and $320\times 320$ pixels, finding better performance with higher resolution.
For MobileNetV3-Small-0.75, performance progressively improves as the resolution increases up to $768\times 768$, but begins to decline at higher resolutions, as shown in Figure 7. In contrast, both ConvNeXt-Tiny and DenseNet-121 exhibit consistent improvements with increasing resolution. ConvNeXt-Tiny outperforms the other two architectures across all evaluated resolutions. The highest performance is achieved with this architecture at a resolution of $768\times 768$, although it surpasses the best result of MobileNetV3-Small-0.75 by only 1.8 $HM$ points. ConvNeXt-Tiny improves upon MobileNetV3-Small-0.75 at the default resolution used in this work ($384\times 384$) by just 2.26 $HM$ points.
This finding is particularly relevant because increasing image resolution has an strong impact on memory requirements and computational cost for both training and inference. In particular, this affects the memory needed for intermediate computations, gradients, and activations within the neural network, making high-resolution training substantially more demanding. Computational cost also increases sharply, as higher resolutions require more multiply-accumulate operations (MACs) in each layer. In addition, the GPU memory usage grows, limiting batch sizes and potentially slowing training. Model complexity further interacts with image resolution: deeper or wider architectures may struggle to process very high-resolution inputs efficiently without optimization strategies such as mixed precision. Because of these constraints, certain experiments could not be performed. For example, training ConvNeXt-Tiny at $1024\times 1024$ and DenseNet-121 at $768\times 768$ and $1024\times 1024$ was not feasible due to GPU memory limitations and excessive computational cost, highlighting a practical limitation in scaling experiments to very high-resolution images.
@@ 6.6 Architectures
We investigate how connectivity patterns and the number of parameters/operations influence pathology classification performance. This is particularly relevant because evidence from the complete data regime in language modeling [scalinglaws, learners] and computer vision [bit2020, imageNet21Kpretraining] suggests that increasing network size and training data consistently reduces error. For meta-learning, chen2019closer evaluated several convolutional architectures in few-shot multi-class classification on mini-ImageNet and a reduced version of the CUB dataset [WahCUB]. Their findings are inconclusive: while deeper architectures improved performance on CUB, gains on mini-ImageNet were observed only in certain cases.
In this experiment, we study the effect of convolutional and attention-based connectivity patterns using popular vision architectures. For both types, we examine efficient architectures with relatively few parameters and operations, as well as larger networks. For efficient convolutional architectures, we focus on ConvNeXt-Atto [convnextv2] and lightweight versions of MobileNet [mobilenets], while for Transformer-based models, we use MobileViTV2-1.0 [mobilevit]. For larger convolutional architectures, we experiment with DenseNet-121, DenseNet-161 [densenet], and ConvNeXt-Tiny, and for Transformer-based models, we consider MobileViTV2-2.0 [mobilevit]. Table 8 summarizes the results, comparing the performance of these architectures along with their number of parameters and operations.
@@
Among the larger architectures, ConvNeXt-Tiny achieves the highest performance, reaching 85.22 $HM$ points. Notably, it outperforms DenseNet-161 by 1.32 $HM$ points, an architecture previously shown to be effective for medical image analysis problems [chestxrayresolutions, learningtodiagnose, pneumoniadetection, limitscohen].
@@
Interestingly, across both efficient and larger architectures, convolutional models outperform their Transformer-based counterparts. This trend is illustrated in Figure 8, which depicts the relationship between model performance and computational efficiency for the evaluated architectures.
@@
Efficient neural network architectures reduce computational requirements without significantly compromising performance, offering several practical benefits. These architectures can run on devices with limited hardware resources, lowering costs and expanding accessibility. Moreover, their efficiency could enable scalable deployment and seamless integration into existing medical infrastructures, including those in remote or resource-constrained regions.
6.7 Hyperparameter analysis
We assess the impact of different hyperparameter configurations on method performance by varying selected hyperparameters and evaluating the resulting classification outcomes. For the BatchBased method, two hyperparameters are tuned during the adaptation phase. The first is the Meta-tst learning rate $lr_{head}$ used to update the head parameters $\varphi$. The second is the proportion of Meta-tst $ptc_{trn}$ examples incorporated into the training steps of each episode $E_{meta-tst}$. In the case of ProtoNet-ML, we investigate two types of encoding layer across different output sizes: a fully connected layer and an average pooling layer. The results of these experiments are shown in Table 9.
For BatchBased, the lower learning rate of $0.005$ consistently yields the best results, regardless of the chosen $ptc_{trn}$ value. Among the three evaluated configurations of $ptc_{trn}$, the highest performance is achieved with a value of $0.5$. In contrast, ProtoNet-ML obtains its best performance when using an average pooling layer with an output size of $128$. Overall, BatchBased shows low sensitivity to hyperparameter variations, maintaining consistent performance across configurations and thus underscoring its robustness.
Table 9: Comparison of model performance across BatchBased and ProtoNet-ML hyperparameter configurations.
[…]
6.8 Visualization of model predictions
We visualize a set of model predictions to further examine its behavior qualitatively. Figure 9 presents selected chest X-ray examples along with their corresponding predictions across the four datasets that comprise MetaChest. The examples are arranged from left to right, progressing from correctly classified cases to those exhibiting substantial errors. For example, in the last row, the image in column (a) shows a PadChest X-ray for which the model correctly predicts all four seen classes as well as the unseen class. In contrast, the image in column (d) of the same row illustrates a case in which the model correctly identifies three categories but misclassifies two pathologies, one seen and one unseen, both in red.
Figure 9: X-ray examples with predicted labels for the 5-way, 1-unseen classification task with 30-shots per class. Each row shows four examples per dataset in MetaChest. Labels below each chest X-ray are annotated with $S$ and $U$ for seen and unseen classes, respectively. Numbers indicate predicted class probabilities; incorrect predictions are shown in red.
@@ 7 Conclusions and future work
The adoption of the generalized few-shot learning paradigm aligns more closely with the clinical presentation of pathologies in chest X-rays than the standard few-shot classification paradigm. This task formulation is particularly well-suited to specialized medical contexts and useful for a variety of scenarios within the healthcare domain. In addition, the proposed multi-label episode generation algorithm enables the creation of complex classification tasks, further broadening its applicability to real-world medical settings. Interestingly, our results show that BatchBased is an effective classification method in few-shot scenarios, despite being based on standard transfer learning and not specifically designed for few-shot learning. We also observed that increasing the number of classes per episode ($n$-way) and the number of training examples per class ($k_{trn}$) improves model performance by enhancing task robustness. With respect to image resolution, we found that using higher resolutions than those commonly applied in natural image tasks leads to better classification performance. This improvement is likely due to the fine-grained nature of pathology classification, where abnormal patterns are subtle and can be overlooked at lower resolutions. However, this performance improvement comes at the cost of higher computational demands and longer training and inference times. In contrast, our results show that efficient architectures can achieve performance comparable to larger models while substantially reducing computational overhead. This is particularly advantageous in resource-constrained environments, such as remote areas or small hospitals, where these architectures strike a balance between performance, computational efficiency, and practical deployability.
As future work, we envision four main research directions. First, leveraging Vision Foundation Models as a starting point for pathology classification. Pre-trained on large-scale datasets, these models could provide richer and more generalizable feature representations, thereby enhancing classification performance. Second, developing multimodal classification models that integrate complementary information from radiology reports, such as radiologist notes and clinical records. Incorporating this additional contextual information could enrich the diagnostic process and improve overall performance. Third, analyzing the behavior of ProtoNet-ML under different distance and activation functions. Finally, conducting a comparative evaluation between model predictions and expert radiologist assessments. Such a study would enable clinical validation of the results and provide a more accurate and reliable measure of the model’s effectiveness in clinical settings.